OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 4, –2, –8, … ?

OpenStudy (anonymous):

To find the value of nth term in a series, we use the following formula\[a_n = a_1 + (n-1) d\] where a1 is the value of the first term, d is the common difference, and n is the number of terms. Here we want to solve for the 20th term. This means n = 20. The above formula becomes \[a_{20} = 4 + (20 -1)(-6)\] lilg132. This is an arithmetic sequence, which means each term differs by a common difference. The series you gave is a geometric series, where each term differens by a common multiple.

OpenStudy (anonymous):

the rule is -6

OpenStudy (anonymous):

Wait so what is the 20th term?

OpenStudy (anonymous):

-8 - 6 = -14, -14 - 6 = -20 -20 - 6 = -26 etc.

OpenStudy (lilg132):

was it -110?

OpenStudy (anonymous):

Use my formula I gave you. 20th term = 4 + 19*(-6)

OpenStudy (anonymous):

-32, -38, -44, -50 that's 10

OpenStudy (anonymous):

@lig u r right

OpenStudy (anonymous):

so -110?

OpenStudy (anonymous):

Yes, -110 is the correct answer.

OpenStudy (anonymous):

Thank you all so much.